OpenStudy (anonymous):

HELP PLEASE Which points are 24 units from the point located at –51 on a number line? A. –75 and 75 B. –75 and –27 C. –27 and 27 D. –27 and 75

OpenStudy (anonymous):

HELP PLEase

OpenStudy (ranga):

You add 24 to -51 to get the point on the right You subtract 24 from -51 to get the point on the left.

OpenStudy (anonymous):

is it C

OpenStudy (ranga):

No. -51 + 24 = -27 -51 - 24 = -75

OpenStudy (anonymous):

SO it is B

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

thank you very much i have another question can you help me

OpenStudy (ranga):

yw. I can try.

OpenStudy (anonymous):

HERE it is Which integer is closest to 0 on the number line? A. –12 B. –8 C. 10 D. 14

OpenStudy (ranga):

Which one do you think it is? Find the distance from 0 by subtracting 0 from each point and take the absolute value. That is, drop the negative sign and compare to see which one is the lowest number. That will be the closest.

OpenStudy (anonymous):

can you give me the answer real quick i have to go

OpenStudy (anonymous):

please